OpenStudy (anonymous):

each letter above the line represents a digit that has difference of one form the digit represented by the same letter below the line. 1, 2, 4, 6, 7, 8. what digit or digits are presented by each letter? A D A E + F C D E A D E E ______ E F D C F

hartnn (hartnn):

i got E,its 2 and so F is 6....working on others

hartnn (hartnn):

can u verify the question please,getting A as 9....which is not in the list 1,2,4,6,7,8

OpenStudy (anonymous):

no, all the the numbers A, C, D, E, F is in the list (1,2,4,6,7,8)

hartnn (hartnn):

since E is carried out,it can be either 1 or 2,if E=1,F=3,which is not in list......so E=2....now (carry)+A+F+A=E*10+F..which gives (carry)+2A=20,since carry can either be 0,1,2...here 0,1 is eliminated,so carry has to be 2.....so I got A=9.....which is still not in the list....

hartnn (hartnn):

since E is carried out,it can be either 1 or 2,if E=1,F=3,which is not in list......so E=2....now (carry)+A+F+A=E*10+F..which gives (carry)+2A=20,since carry can either be 0,1,2...here 0,1 is eliminated,so carry has to be 2.....so I got A=9.....which is still not in the list....